A 6.62 L sample of oxygen gas is collected at 46 C. Calculate the volume of the gas when the temperature is at standard conditions (STP). Assume constant pressure.
1
Expert's answer
2019-02-04T09:37:06-0500

V/T = V1/T1

V1= V*T1 / T

V1 = 6.62*273/ 319

V1= 5.66 L